# On tori periods of Weil representations of unitary groups

### v1

@@ 2.4. Admissible embeddings
Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $K$. Let $E$ be an étale algebra of rank $n$ over $F$. In this section, we define the notion of an admissible embedding $i\colon L_{E}^{1}\rightarrow\operatorname{U}(\mathrm{V})$. We classify the set of admissible embeddings modulo conjugation by elements of $\operatorname{U}(\mathrm{V})$. These results are needed for Section 5. We say that an embedding $i\colon L_{E}^{1}\rightarrow\operatorname{U}(\mathrm{V})$ is admissible if there exists $\lambda\in E^{\times}$ and an isomorphism of hermitian spaces $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$, such that for any $x\in L_{E}^{1}$,
@@
for every $x\in L_{E}^{1}$, if and only if there exists $y\in L_{E}^{\times}$, such that $\lambda=N_{L_{E}/E}\left(y\right)\lambda^{\prime}$ and $r_{1}=r_{2}\circ\mathrm{m}_{y}$. Indeed, if $\lambda=\lambda^{\prime}\cdot N_{L_{E}/E}\left(y\right)$ for $y\in L_{E}^{\times}$ and $r_{1}=r_{2}\circ\mathrm{m}_{y}$, it is easy to check that (3) holds. On the other hand if (3) holds, we have from Proposition 2.5 below that $r_{2}^{-1}\circ r_{1}\colon L_{E}\to L_{E}$ is an isomorphism of $L_{E}$-algebras, and therefore there exists $y\in L_{E}$ such that $r_{2}^{-1}\circ r_{1}=\mathrm{m}_{y}$. Since $r_{2}^{-1}\circ r_{1}\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}_{E,\lambda^{\prime}}$ is an isomorphism of hermitian spaces, we must have that for every $x_{1},x_{2}\in L_{E}$,
@@
$i_{r^{\prime}}\left(x\right)=r^{\prime}\circ{r}^{-1}\circ i\left(x\right)\circ r\circ{r^{\prime}}^{-1},$
@@
is an $L_{E}$-linear map. Suppose that there exists $h\in\operatorname{U}(\mathrm{V})$ such that ${r^{-1}\circ h\circ r^{\prime}}$ is an $L_{E}$-linear map. Then $h\circ r^{\prime}=r\circ\mathrm{m}_{y}$ for some $y\in L_{E}^{\times}$. For any $x_{1},x_{2}\in L_{E}$, we have
@@
as required. We move to prove the other direction. Suppose that $\lambda^{\prime}=N_{L_{E}/E}\left(y\right)\lambda$, where $y\in L_{E}^{\times}$. Define $h=r\circ\mathrm{m}_{y}\circ{r^{\prime}}^{-1}.$ Then $r^{-1}\circ h\circ r^{\prime}=\mathrm{m}_{y}$ is an $L_{E}$-linear map. We check that the element $h$ we have constructed lies in the unitary group $\operatorname{U}(\mathrm{V})$. Repeating similar steps to before, we have
@@
Here, $\bar{F}$ is the algebraic closure of $F$, and we refer to Section 2.5 for the definition of $\operatorname{U}(\mathrm{V})\left(\bar{F}\right)$. We have that the set of admissible embeddings $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is a stable conjugacy class. We will show this only in the case where $K/F$ is a quadratic field extension, as the other case is simpler. To see this, first recall that $\operatorname{U}(\mathrm{V})\left(\bar{F}\right)$ is naturally isomorphic to $\operatorname{Res}_{\bar{F}/F}\operatorname{GL}\left(\mathrm{V}\otimes_{K}\bar{F}\right)$, that is, the group of invertible $\bar{F}$-linear maps $\mathrm{V}\otimes_{K}\bar{F}\to\mathrm{V}\otimes_{K}\bar{F}.$ If $i_{1}$ and $i_{2}$ are admissible embeddings corresponding to the data $r_{1}\colon\mathrm{V}_{E,\lambda_{1}}\to\mathrm{V}$ and $r_{2}\colon\mathrm{V}_{E,\lambda_{2}}\to\mathrm{V}$ with $\lambda_{1},\lambda_{2}\in E^{\times}$, then for any $x\in L_{E}^{1}$, we have $g\circ i_{1}\left(x\right)\circ g^{-1}=i_{2}\left(x\right),$ where $g=r_{2}\circ r_{1}^{-1}\in\operatorname{Res}_{K/F}\operatorname{GL}\left(\mathrm{V}\right)$. Hence, $i_{1}$ and $i_{2}$ are stably conjugate. On the other hand, if $i_{1}\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is an admissible embedding, corresponding to the data $\lambda\in E^{\times}$ and $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$, and $i_{2}\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is an embedding, such that $i_{1}$ and $i_{2}$ are stably conjugate, then there exists an invertible $\bar{F}$-linear map $g\colon\mathrm{V}\otimes_{K}\bar{F}\to\mathrm{V}\otimes_{K}\bar{F}$ such that $g\circ\left(i_{1}\left(x\right)\otimes\operatorname{id}_{\overline{F}}\right)\circ g^{-1}=i_{2}\left(x\right)\otimes\operatorname{id}_{\overline{F}}$ for any $x\in L_{E}^{1}$. Since $i_{2}\left(L_{E}^{1}\right)$ is a maximal torus in $\operatorname{U}(\mathrm{V})$, there exists an étale algebra $E^{\prime}$ of rank $n$ over $F$, a $\lambda^{\prime}\in\left(E^{\prime}\right)^{\times}$, and an isomorphism of hermitian spaces $r^{\prime}\colon\mathrm{V}_{E^{\prime},{\lambda^{\prime}}}\to\mathrm{V}$, such that the image of the map $L_{E^{\prime}}^{1}\to\operatorname{U}(\mathrm{V})$ given by $x\mapsto r^{\prime}\circ\mathrm{m}_{x}\circ\left(r^{\prime}\right)^{-1}$ is $i_{2}\left(L_{E}^{1}\right)$. Consider the map $T=\left(r^{-1}\otimes\operatorname{id}_{\bar{F}}\right)\circ g^{-1}\circ\left(r^{\prime}\otimes\operatorname{id}_{\overline{F}}\right)\colon L_{E^{\prime}}\otimes_{K}\overline{F}\to L_{E}\otimes_{K}\overline{F}$. Since we have for any $x\in L_{E}^{1}$, $T\circ\left(\mathrm{m}_{x}\otimes\operatorname{id}_{\bar{F}}\right)=\left(\mathrm{m}_{\tau\left(x\right)}\otimes\operatorname{id}_{\bar{F}}\right)\circ T$ for some $\tau\left(x\right)\in L_{E^{\prime}}^{1}$, by Proposition A.2 for any $x\in L_{E}$, there exists a unique $\tau\left(x\right)\in L_{E^{\prime}}$, such that $T\circ\left(\mathrm{m}_{\tau\left(x\right)}\otimes\operatorname{id}_{\bar{F}}\right)=\left(\mathrm{m}_{x}\otimes\operatorname{id}_{\bar{F}}\right)\circ T$. It follows that $\tau\colon L_{E}\to L_{E^{\prime}}$ is an isomorphism of étale algebras over $K$, that is, it is an isomorphism of rings that fixes $K$. We have that for any $x\in L_{E}^{1}$,
@@
$i_{r_{2}}\left(x\right)=r_{2}\circ{r}^{-1}\circ i_{r}\left(x\right)\circ r\circ{r_{2}}^{-1},$
and we have that $r\circ{r_{2}}^{-1}\in\operatorname{U}(\mathrm{V})$ as a composition of two isometries. We show that $\lambda\mapsto c_{\lambda}$ is a bijection as in the theorem. The map is injective: given $\lambda,\lambda^{\prime}\in E^{\times}$, and isomorphisms $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$ and ${r^{\prime}\colon\mathrm{V}_{E,\lambda^{\prime}}\rightarrow\mathrm{V}}$, we have for any $x\in L_{E}^{1}$,
$i_{r^{\prime}}\left(x\right)=r^{\prime}\circ{r}^{-1}\circ i_{r}\left(x\right)\circ r\circ{r^{\prime}}^{-1}.$
It follows from Lemma 2.6 that the embeddings $i_{r}$ and $i_{r^{\prime}}$ are conjugate if and only if $\left(\lambda^{\prime}\right)^{-1}\lambda\in N_{L_{E}/E}\left(L_{E}^{\times}\right)$. The map is surjective: let $i\colon L_{E}^{1}\rightarrow\operatorname{U}(\mathrm{V})$ be an admissible embedding corresponding to the data ${r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}}$. We have that for any $x\in L_{E}^{1}$,
@@
Let $\lambda,\lambda^{\prime}\in E^{\times}$. If $F$ is a non-archimedean local field then, as in the previous remark, we have that the hermitian spaces $\mathrm{V}_{E,\lambda}$ and $\mathrm{V}_{E,\lambda^{\prime}}$ are isomorphic if and only if $N_{E/F}\left(\lambda^{-1}\lambda^{\prime}\right)\in N_{K/F}\left(K^{\times}\right)$. If $F$ is an archimedean local field, then the hermitian spaces $\mathrm{V}_{E,\lambda}$ and $\mathrm{V}_{E,\lambda^{\prime}}$ are isomorphic if and only if the number of non-trivial components of $\lambda$ and $\lambda^{\prime}$ as elements of $E^{\times}/N_{L_{E}/E}\left(L_{E}^{\times}\right)$ is the same.
@@ 2.5. Functoriality for admissible embeddings
In this section, we explain how an admissible embedding of a torus $i\colon L_{E}^{1}\to\mathrm{T}\subset\operatorname{U}(\mathrm{V})$, gives rise to a family of embeddings $i\left(R\right)\colon L_{E}^{1}\left(R\right)\to\operatorname{U}(\mathrm{V})\left(R\right)$ for any $F$-algebra $R$. Let $R$ be a commutative $F$-algebra, and consider the ring $L_{R}=K\otimes_{F}R$, equipped with the involution $\sigma\colon L_{R}\to L_{R}$ defined on pure tensors by
@@
where $k\in K$ and $h\in R$. Henceforth, we will write $x^{c}$ instead of $\sigma\left(x\right)$ for $x\in L_{R}$. As before, one can define the notion of an $\epsilon$-hermitian space over $L_{R}$, as in Section 2.1 by replacing $K$ with $L_{R}$ in the definitions. Let $\left(\mathrm{V},\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}}\right)$ be an $\epsilon$-hermitian space over $K$. Then the space $(\mathrm{V}\left(R\right),\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}\left(R\right)})$ is an $\epsilon$-hermitian space over $L_{R}$, where $\mathrm{V}\left(R\right)=\mathrm{V}\otimes_{F}R$ and $\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}\left(R\right)}$ is defined on pure tensors by
@@
where $v_{1},v_{2}\in\mathrm{V}$ and $h_{1},h_{2}\in R$. We denote by $\operatorname{U}(\mathrm{V})\left(R\right)$ the group consisting of invertible $L_{R}$-linear maps that preserve the form $\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}\left(R\right)}$. Suppose that $E$ is an étale algebra over $F$ of degree $n=\dim\mathrm{V}$, and that $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is an admissible embedding corresponding to the data $\lambda\in E^{\times}$ and $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$. Since $r$ is an isomorphism of $\epsilon$-hermitian spaces over $K$, we have that the map $r\left(R\right)=r\otimes\operatorname{id}_{R}\colon\mathrm{V}_{E,\lambda}\left(R\right)\to\mathrm{V}\left(R\right)$ is an isomorphism of $\epsilon$-hermitian spaces over $L_{R}$. We may define an embedding $i\left(R\right)\colon L_{E}^{1}\left(R\right)\to\operatorname{U}(\mathrm{V})\left(R\right)$, corresponding to the data $\lambda\in E^{\times}$ and $r\left(R\right)\colon\mathrm{V}_{E,\lambda}\left(R\right)\to\mathrm{V}\left(R\right)$, by the formula
@@
$L_{E}^{1}\left(R\right)=\left\{x\in\left(L_{E}\otimes_{F}R\right)^{\times}\mid x\cdot x^{c}=1\right\},$
@@
where $y\in L_{E}$ and $h\in R$. Note that if $k\in K^{1}$, we always have that $i\left(R\right)\left(k\otimes 1\right)=\mathrm{m}_{k}\otimes\operatorname{id}_{R}$, where $\mathrm{m}_{k}\colon\mathrm{V}\to\mathrm{V}$ is the map $\mathrm{m}_{k}v=kv$, i.e., the multiplication by the scalar $k$ map. In the special case where $\mathrm{V}$ is one-dimensional, and hence $E=F$, we get that $i\left(R\right)\colon L_{E}^{1}\left(R\right)\to\operatorname{U}(\mathrm{V})\left(R\right)$ is an isomorphism for every $R$. In the sequel, given an admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$, we will often write $i$ for $i\left(R\right)$, especially when $F=\mathbf{F}$ is a number field and $R=\mathbb{A}_{\mathbf{F}}$ is its ring of adeles.

### v2

@@ 2.4. Admissible embeddings
Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $K$. Let $E$ be an étale algebra of rank $n$ over $F$. In this section, we define the notion of an admissible embedding $i\colon L_{E}^{1}\rightarrow\operatorname{U}(\mathrm{V})$. We classify the set of admissible embeddings modulo conjugation by elements of $\operatorname{U}(\mathrm{V})$. These results are needed for Section 5.
We say that an embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is admissible if there exists $\lambda\in E^{\times}$ and an isomorphism of hermitian spaces $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$, such that for any $x\in L_{E}^{1}$,
@@
for every $x\in L_{E}^{1}$, if and only if there exists $y\in L_{E}^{\times}$, such that $\lambda=N_{L_{E}/E}\left(y\right)\lambda^{\prime}$ and $r_{1}=r_{2}\circ\mathrm{m}_{y}$.
Indeed, if $\lambda=\lambda^{\prime}\cdot N_{L_{E}/E}\left(y\right)$ for $y\in L_{E}^{\times}$ and $r_{1}=r_{2}\circ\mathrm{m}_{y}$, it is easy to check that (3) holds. On the other hand if (3) holds, we have from Proposition 2.5 below that $r_{2}^{-1}\circ r_{1}\colon L_{E}\to L_{E}$ is an isomorphism of $L_{E}$-algebras, and therefore there exists $y\in L_{E}$ such that $r_{2}^{-1}\circ r_{1}=\mathrm{m}_{y}$. Since $r_{2}^{-1}\circ r_{1}\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}_{E,\lambda^{\prime}}$ is an isomorphism of hermitian spaces, we must have that for every $x_{1},x_{2}\in L_{E}$,
@@
$i_{r^{\prime}}\left(x\right)=\left(r\circ{r^{\prime}}^{-1}\right)^{-1}\circ i\left(x\right)\circ\left(r\circ{r^{\prime}}^{-1}\right),$
@@
is an $L_{E}$-linear map.
Suppose that there exists $h\in\operatorname{U}(\mathrm{V})$ such that ${r^{-1}\circ h\circ r^{\prime}}$ is an $L_{E}$-linear map. Then $h\circ r^{\prime}=r\circ\mathrm{m}_{y}$ for some $y\in L_{E}^{\times}$. For any $x_{1},x_{2}\in L_{E}$, we have
@@
as required.
We move to prove the other direction. Suppose that $\lambda^{\prime}=N_{L_{E}/E}\left(y\right)\lambda$, where $y\in L_{E}^{\times}$. Define $h=r\circ\mathrm{m}_{y}\circ{r^{\prime}}^{-1}.$ Then $r^{-1}\circ h\circ r^{\prime}=\mathrm{m}_{y}$ is an $L_{E}$-linear map. We check that the element $h$ we have constructed lies in the unitary group $\operatorname{U}(\mathrm{V})$. Repeating similar steps to before, we have
@@
Here, $\bar{F}$ is an algebraic closure of $F$, and we refer to Section 2.5 for the definition of $\operatorname{U}(\mathrm{V})\left(\bar{F}\right)$. We have that the set of admissible embeddings $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is a stable conjugacy class. We will show this only in the case where $K/F$ is a quadratic field extension, as the other case is simpler. In this case, we choose $\bar{F}$ to be an algebraic closure of $K$. We have that $\operatorname{U}(\mathrm{V})\left(\bar{F}\right)$ is naturally isomorphic to $\operatorname{GL}_{\bar{F}}\left(\mathrm{V}\otimes_{K}\bar{F}\right)$, the group of invertible $\bar{F}$-linear maps $\mathrm{V}\otimes_{K}\bar{F}\to\mathrm{V}\otimes_{K}\bar{F}.$ If $i_{1}$ and $i_{2}$ are admissible embeddings corresponding to the data $r_{1}\colon\mathrm{V}_{E,\lambda_{1}}\to\mathrm{V}$ and $r_{2}\colon\mathrm{V}_{E,\lambda_{2}}\to\mathrm{V}$ with $\lambda_{1},\lambda_{2}\in E^{\times}$, then for any $x\in L_{E}^{1}$, we have $g\circ i_{1}\left(x\right)\circ g^{-1}=i_{2}\left(x\right),$ where $g=r_{2}\circ r_{1}^{-1}\in\operatorname{GL}_{K}\left(\mathrm{V}\right)$ (where $\operatorname{GL}_{K}\left(\mathrm{V}\right)$ is the group of invertible $K$-linear maps $\mathrm{V}\to\mathrm{V}$). Hence, $i_{1}$ and $i_{2}$ are stably conjugate.
On the other hand, if $i_{1}\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is an admissible embedding, corresponding to the data $\lambda\in E^{\times}$ and $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$, and $i_{2}\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is an embedding, such that $i_{1}$ and $i_{2}$ are stably conjugate, then there exists an invertible $\bar{F}$-linear map $g\colon\mathrm{V}\otimes_{K}\bar{F}\to\mathrm{V}\otimes_{K}\bar{F}$ such that $g\circ\left(i_{1}\left(x\right)\otimes\operatorname{id}_{\bar{F}}\right)\circ g^{-1}=i_{2}\left(x\right)\otimes\operatorname{id}_{\bar{F}}$ for any $x\in L_{E}^{1}$. Since $i_{2}\left(L_{E}^{1}\right)$ is a maximal torus in $\operatorname{U}(\mathrm{V})$, there exists an étale algebra $E^{\prime}$ of rank $n$ over $F$, a $\lambda^{\prime}\in\left(E^{\prime}\right)^{\times}$, and an isomorphism of hermitian spaces $r^{\prime}\colon\mathrm{V}_{E^{\prime},{\lambda^{\prime}}}\to\mathrm{V}$, such that the image of the map $L_{E^{\prime}}^{1}\to\operatorname{U}(\mathrm{V})$ given by $x\mapsto r^{\prime}\circ\mathrm{m}_{x}\circ\left(r^{\prime}\right)^{-1}$ is $i_{2}\left(L_{E}^{1}\right)$. Consider the map $T=\left(r^{-1}\otimes\operatorname{id}_{\bar{F}}\right)\circ g^{-1}\circ\left(r^{\prime}\otimes\operatorname{id}_{\bar{F}}\right)\colon L_{E^{\prime}}\otimes_{K}\bar{F}\to L_{E}\otimes_{K}\bar{F}$. Since we have for any $x\in L_{E}^{1}$, $T\circ\left(\mathrm{m}_{x}\otimes\operatorname{id}_{\bar{F}}\right)=\left(\mathrm{m}_{\tau\left(x\right)}\otimes\operatorname{id}_{\bar{F}}\right)\circ T$ for some $\tau\left(x\right)\in L_{E^{\prime}}^{1}$, by Proposition A.2 for any $x\in L_{E}$, there exists a unique $\tau\left(x\right)\in L_{E^{\prime}}$, such that $T\circ\left(\mathrm{m}_{\tau\left(x\right)}\otimes\operatorname{id}_{\bar{F}}\right)=\left(\mathrm{m}_{x}\otimes\operatorname{id}_{\bar{F}}\right)\circ T$. It follows that $\tau\colon L_{E}\to L_{E^{\prime}}$ is an isomorphism of étale algebras over $K$, that is, it is an isomorphism of rings that fixes $K$. We have that for any $x\in L_{E}^{1}$,
@@
$i_{r_{2}}\left(x\right)=\left(r\circ r_{2}^{-1}\right)^{-1}\circ i_{r}\left(x\right)\circ r\circ{r_{2}}^{-1},$
and we have that $r\circ{r_{2}}^{-1}\in\operatorname{U}(\mathrm{V})$ as a composition of two isometries.
We show that $\lambda\mapsto c_{\lambda}$ is a bijection as in the theorem.
The map is injective: given $\lambda,\lambda^{\prime}\in E^{\times}$, and isomorphisms $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$ and ${r^{\prime}\colon\mathrm{V}_{E,\lambda^{\prime}}\rightarrow\mathrm{V}}$, we have for any $x\in L_{E}^{1}$,
$i_{r^{\prime}}\left(x\right)=\left({r}\circ r^{\prime-1}\right)^{-1}\circ i_{r}\left(x\right)\circ r\circ{r^{\prime}}^{-1}.$
It follows from Lemma 2.6 that the embeddings $i_{r}$ and $i_{r^{\prime}}$ are conjugate if and only if $\left(\lambda^{\prime}\right)^{-1}\lambda\in N_{L_{E}/E}\left(L_{E}^{\times}\right)$.
The map is surjective: let $i\colon L_{E}^{1}\rightarrow\operatorname{U}(\mathrm{V})$ be an admissible embedding corresponding to the data ${r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}}$. We have that for any $x\in L_{E}^{1}$,
@@
Let $\lambda,\lambda^{\prime}\in E^{\times}$. If $F$ is a non-archimedean local field then, as in the previous remark, we have that the hermitian spaces $\mathrm{V}_{E,\lambda}$ and $\mathrm{V}_{E,\lambda^{\prime}}$ are isomorphic if and only if $N_{E/F}\left(\lambda^{-1}\lambda^{\prime}\right)\in N_{K/F}\left(K^{\times}\right)$.
If $F$ is an archimedean local field, then the hermitian spaces $\mathrm{V}_{E,\lambda}$ and $\mathrm{V}_{E,\lambda^{\prime}}$ are isomorphic if and only if the number of non-trivial components of $\lambda$ and $\lambda^{\prime}$ as elements of $E^{\times}/N_{L_{E}/E}\left(L_{E}^{\times}\right)$ is the same.
@@ 2.5. Functoriality for admissible embeddings
In this section, we explain how an admissible embedding of a torus $i\colon L_{E}^{1}\to\mathrm{T}\subset\operatorname{U}(\mathrm{V})$, gives rise to a family of embeddings $i\left(R\right)\colon L_{E}^{1}\left(R\right)\to\operatorname{U}(\mathrm{V})\left(R\right)$ for any $F$-algebra $R$.
Let $R$ be a commutative $F$-algebra, and consider the ring $L_{R}=K\otimes_{F}R$, equipped with the involution $\sigma\colon L_{R}\to L_{R}$ defined on pure tensors by
@@
where $k\in K$ and $h\in R$. Henceforth, we will write $x^{c}$ instead of $\sigma\left(x\right)$ for $x\in L_{R}$.
As before, one can define the notion of an $\epsilon$-hermitian space over $L_{R}$, as in Section 2.1 by replacing $K$ with $L_{R}$ in the definitions.
Let $\left(\mathrm{V},\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}}\right)$ be an $\epsilon$-hermitian space over $K$. Then the space $(\mathrm{V}\left(R\right),\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}\left(R\right)})$ is an $\epsilon$-hermitian space over $L_{R}$, where $\mathrm{V}\left(R\right)=\mathrm{V}\otimes_{F}R$ and $\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}\left(R\right)}$ is defined on pure tensors by
@@
where $v_{1},v_{2}\in\mathrm{V}$ and $h_{1},h_{2}\in R$. We denote by $\operatorname{U}(\mathrm{V})\left(R\right)$ the group consisting of invertible $L_{R}$-linear maps that preserve the form $\left\langle\cdot,\cdot\right\rangle_{\mathrm{V}\left(R\right)}$. Suppose that $E$ is an étale algebra over $F$ of degree $n=\dim\mathrm{V}$, and that $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$ is an admissible embedding corresponding to the data $\lambda\in E^{\times}$ and $r\colon\mathrm{V}_{E,\lambda}\to\mathrm{V}$. Since $r$ is an isomorphism of $\epsilon$-hermitian spaces over $K$, we have that the map $r\left(R\right)=r\otimes\operatorname{id}_{R}\colon\mathrm{V}_{E,\lambda}\left(R\right)\to\mathrm{V}\left(R\right)$ is an isomorphism of $\epsilon$-hermitian spaces over $L_{R}$.
We may define an embedding $i\left(R\right)\colon\operatorname{Res}_{E/F}L_{E}^{1}\left(R\right)\to\operatorname{U}(\mathrm{V})\left(R\right)$, corresponding to the data $\lambda\in E^{\times}$ and $r\left(R\right)\colon\mathrm{V}_{E,\lambda}\left(R\right)\to\mathrm{V}\left(R\right)$, by the formula
@@
$\operatorname{Res}_{E/F}L_{E}^{1}\left(R\right)=\left\{x\in\left(L_{E}\otimes_{F}R\right)^{\times}\mid x\cdot x^{c}=1\right\},$
@@
where $y\in L_{E}$ and $h\in R$.
Note that if $k\in K^{1}$, we always have that $i\left(R\right)\left(k\otimes 1\right)=\mathrm{m}_{k}\otimes\operatorname{id}_{R}$, where $\mathrm{m}_{k}\colon\mathrm{V}\to\mathrm{V}$ is the map $\mathrm{m}_{k}v=kv$, i.e., the multiplication by the scalar $k$ map. In the special case where $\mathrm{V}$ is one-dimensional, and hence $E=F$, we get that $i\left(R\right)\colon L_{E}^{1}\left(R\right)\to\operatorname{U}(\mathrm{V})\left(R\right)$ is an isomorphism for every $R$.
In the sequel, given an admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{V})$, we will often write $i$ for $i\left(R\right)$, especially when $F=\mathbf{F}$ is a number field and $R=\mathbb{A}_{\mathbf{F}}$ is its ring of adeles.